OpenStudy (anonymous):

give tan(theta) =4/3 and sin(-4/5) find cos(theta) 1/3 -1/3 3/5 -3/5 -sqrt3/5

ganeshie8 (ganeshie8):

hint : \(\large \tan \theta = \frac{\sin \theta}{\cos \theta} \)

OpenStudy (anonymous):

4/3*-4/5=cosθ

OpenStudy (anonymous):

i got 8/15? but i know thats wrong

OpenStudy (anonymous):

i got 8/15 wheni tried harsh123 solution and i got no solution for gabeshie8 solution

OpenStudy (anonymous):

*ganeshie8

ganeshie8 (ganeshie8):

\(\large \tan \theta = \frac{\sin \theta}{\cos \theta}\) \(\large \frac{4}{3} = \frac{\frac{-4}{5}}{\cos \theta}\) cross multiply \(\large 4 \cos \theta= 3\times \frac{-4}{5}\) divide 4 both sides \(\large \cos \theta= \frac{-3}{5}\)

OpenStudy (anonymous):

oh thankyou

ganeshie8 (ganeshie8):

np :)